A bus travelling first one third dist. at speed 10 km/h and next one fourth at ...

Grade:12th

1 Answers

Arun
6 years ago
the total distance covered be x
thus
time takes to travel 1/3x = t1 = x/30
 
time taken to travel 1/4x = t2 = x/80
 
time taken to travel the rest = t3 = 5x/480 = x/96
 
thus average speed = total distnace/total time taken = 1/(1/30 + 1/80 + 1/96) kmph
Regards
Arun (askIITians forum expert)